# Fast Benchmarking of Asynchronous Multi-Fidelity Optimization on Zero-Cost Benchmarks

### v1

@@ 5  Experiments on Zero-Cost Benchmarks Using Various Open-Sourced HPO Tools
The aim of this section is to show that: (1) our wrapper is applicable to diverse HPO libraries and HPO benchmarks, and that (2) ranking of algorithms varies under benchmarking of parallel setups, making such evaluations necessary. We use random search and TPE (Bergstra et al.,, 2011; Watanabe, 2023b, ) from Optuna (Akiba et al.,, 2019), random forest-based Bayesian optimization (via the MFFacade) from SMAC3 (Lindauer et al.,, 2022), DEHB (Awad et al.,, 2021), HyperBand (Li et al.,, 2017) and BOHB (Falkner et al.,, 2018) from HpBandSter, NePS 555 It was under development when we used it and the package is available at https://github.com/automl/neps/. , and HEBO (Cowen-Rivers et al.,, 2022) as optimizers. For more details, see Appendix B. Optuna uses multithreading, SMAC3 and DEHB use dask, HpBandSter uses file server-based synchronization, NePS uses file system-based synchronization, and HEBO uses the ask-and-tell interface. In the experiments, we used these optimizers with our wrapper to optimize the MLP benchmark in HPOBench (Eggensperger et al.,, 2021), HPOLib (Klein and Hutter,, 2019), JAHS-Bench-201 (Bansal et al.,, 2022), LCBench (Zimmer et al.,, 2021) in YAHPOBench (Pfisterer et al.,, 2022), and two multi-fidelity benchmark functions proposed by Kandasamy et al., (2017). See Appendix A for more details. We used the number of parallel workers $P\in\{1,2,4,8\}$ over 30 different random seeds for each and $\eta=3$ for HyperBand-based methods, i.e. the default value of a control parameter of HyperBand that determines the proportion of HP configurations discarded in each round of successive halving (Jamieson and Talwalkar,, 2016). The budget for each optimization was fixed to 200 full evaluations and this leads to 450 function calls for HyperBand-based methods with $\eta=3$. Note that random search and HyperBand used 10 times more budget, i.e. 2000 full evaluations, compared to the others. All the experiments were performed on bwForCluster NEMO, which has 20 cores of Intel(R) Xeon(R) CPU E5-2630 v4 on each computational node, and we used 15GB RAM per worker.
@@
Table 1: Actual and simulated runtimes for total experiment runtimes. Act.: total actual runtime and Sim.: total simulated runtime. $\times$ Fast: speedup factor of simulation.
@@ 6  Broader Impact & Limitations
The primary motivation for this paper is to reduce the runtime of simulations for MFO. As shown in Table 1, our experiments would have taken $5.6\times 10^{10}$ seconds $\simeq 1.8\times 10^{3}$ CPU years with the naïve simulation. As a computer consumes about 30Wh even just for waiting and $736~{}\mathrm{g}$ of $\mathrm{CO}_{2}$ is produced per 1kWh, the whole experiment would have produced about $3.5\times 10^{2}~{}\mathrm{t}$ of $\mathrm{CO}_{2}$. It means that our wrapper saved $3.497\times 10^{2}~{}\mathrm{t}$ of $\mathrm{CO}_{2}$ production. Therefore, researchers can also reduce the similar amount of $\mathrm{CO}_{2}$ for each experiment. The main limitation of our current wrapper is the assumption that none of the workers will not die and any additional workers will not be added after the initialization. Besides that, our package cannot be used on Windows OS because fcntl is not supported on Windows.

### v3

@@ 5  Experiments on Zero-Cost Benchmarks Using Various Open-Sourced HPO Tools
The aim of this section is to show that: (1) our wrapper is applicable to diverse HPO libraries and HPO benchmarks, and that (2) ranking of algorithms varies under benchmarking of parallel setups, making such evaluations necessary. We use random search and TPE (Bergstra et al.,, 2011; Watanabe, 2023b, ) from Optuna (Akiba et al.,, 2019), random forest-based Bayesian optimization (via the MFFacade) from SMAC3 (Lindauer et al.,, 2022), DEHB (Awad et al.,, 2021), HyperBand (Li et al.,, 2017) and BOHB (Falkner et al.,, 2018) from HpBandSter, NePS 555 It was under development when we used it and the package is available at https://github.com/automl/neps/. , and HEBO (Cowen-Rivers et al.,, 2022) as optimizers. For more details, see Appendix B. Optuna uses multithreading, SMAC3 and DEHB use dask, HpBandSter uses file server-based synchronization, NePS uses file system-based synchronization, and HEBO uses the ask-and-tell interface. In the experiments, we used these optimizers with our wrapper to optimize the MLP benchmark in HPOBench (Eggensperger et al.,, 2021), HPOLib (Klein and Hutter,, 2019), JAHS-Bench-201 (Bansal et al.,, 2022), LCBench (Zimmer et al.,, 2021) in YAHPOBench (Pfisterer et al.,, 2022), and two multi-fidelity benchmark functions proposed by Kandasamy et al., (2017). See Appendix A for more details. We used the number of parallel workers $P\in\{1,2,4,8\}$ over 30 different random seeds for each and $\eta=3$ for HyperBand-based methods, i.e. the default value of a control parameter of HyperBand that determines the proportion of HP configurations discarded in each round of successive halving (Jamieson and Talwalkar,, 2016). The budget for each optimization was fixed to 200 full evaluations and this leads to 450 function calls for HyperBand-based methods with $\eta=3$. Note that random search and HyperBand used 10 times more budget, i.e. 2000 full evaluations, compared to the others. All the experiments were performed on bwForCluster NEMO, which has 10 cores of Intel(R) Xeon(R) CPU E5-2630 v4 on each computational node, and we used 15GB RAM per worker.
@@
Table 1: The total actual and simulated runtimes over all the experiments. Act.: total actual runtime and Sim.: total simulated runtime. $\times$ Fast: speedup factor of simulation.
@@ 6  Broader Impact & Limitations
The primary motivation for this paper is to reduce the runtime of simulations for MFO. As shown in Table 1, our experiments would have taken $5.6\times 10^{10}$ seconds $\simeq 1.8\times 10^{3}$ CPU years with the naïve simulation. As the TDP of Intel(R) Xeon(R) CPU E5-2630 v4 used in our experiments consumes about 85W and about $350~{}\mathrm{g}$ of $\mathrm{CO}_{2}$ is produced per 1kWh, the whole experiment would have produced about $9.333~{}\mathrm{t}$ of $\mathrm{CO}_{2}$ if we estimate a core of the CPU needs 2W in its idole state. It means that our wrapper saved $9.326~{}\mathrm{t}$ of $\mathrm{CO}_{2}$ production at least. Therefore, researchers can also reduce the similar amount of $\mathrm{CO}_{2}$ for each experiment. The main limitation of our current wrapper is the assumption that none of the workers will not die and any additional workers will not be added after the initialization. Besides that, our package cannot be used on Windows OS because fcntl is not supported on Windows.